Grade 11General Physics

If the equation for the angular displacement of a particle moving on a circular path is given by theta=2t3 + 0.5 where theta is in radians and t in seconds, then the average angular velocity of the particle after 2 seconds from its start is?

Rituraj Singh
12 Years agoGrade 11

3 Answers

Arun Kumar
12 Years ago
Hello Student,
\theta=2t3 + 0.5
\theta =8.5
t=2
so average =4.25rad/s
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
Nidhi
8 Years ago
Theta = 2t^3+0.5 ...(given)...(1)
We know that ,
Omega =d(theta)/dt
therefore, 
Diff. eqn(1) wrt t,
So, 
Omega= 6t^2 
At t=2 ,
Omega = 6×4=24 rad/s
(Theta is displacement, Omega is angular velocity)
Hope this helps
Kushagra Madhukar
5 Years ago
Dear student,
Please find the solution to your problem below.
 
SInce here average angular velocity is asked and not the instantaneous angular velocity, we will have to find the total change in theta in the given 2s instead of differentiating.
Now, at t = 0,  θo = 2(0)3 + 0.5 = 0.5 rad.
at t = 2s, θ2 = 2(2)3 + 0.5 = 8.5 rad.
Average angular velocity, w = 2 – θo)/(2 – 0)
                                        = 8/2
                                        = 4 rad/s
 
Hope this helps.
Thanks and regards,
Kushagra